# Concurrence fill and mode distribution of entanglement in neutrino oscillation

### v1

@@ C. Tangle (or Three-Tangle) and Partial Tangle
Figure 4: Energy dependence of the partial tangles $\tau^{2}_{i|jk}$ in the three-flavor neutrino system for vacuum (dashed) and matter (solid) propagation. The curves correspond to the three bipartitions $\tau^{2}_{\mu e}$ (tuquoise), $\tau^{2}_{e\tau}$ (purple), and $\tau^{2}_{\mu\tau}$ (yellow), and quantify how the residual entanglement associated with each flavor mode is shared among the remaining two modes. Vertical markers indicate representative energies (2.6, 1.7, 1.3, and 0.8 GeV) corresponding to the first, intermediate, and second oscillation maxima. The variation of the partial tangles demonstrates that the distribution of residual entanglement evolves non-uniformly with energy: in vacuum, the structure reflects the coherent interplay among flavor components, whereas matter effects further reshape this pattern by enhancing or suppressing individual bipartitions. Overall, the figure provides a compact view of how tripartite entanglement is dynamically reorganized as the neutrino propagates, offering insight into the quantum structure relevant for long-baseline experiments such as DUNE.
@@
We next examine the partial tangle, $\tau_{\alpha\beta}$, which probes the residual bipartite entanglement between a selected pair $(\alpha,\beta)$ within a correlated tripartite system. Figure 4 displays the energy dependence of the three partial tangles $\{\tau_{\mu e}^{2}\ \text{(cyan)},\ \tau_{\mu\tau}^{2}\ \text{(yellow)},\ \tau_{e\tau}^{2}\ \text{(purple)}\}$ in both vacuum (dashed) and matter (solid) across the entire accessible energy range of DUNE. We observe that each partial tangle lies in the interval $[0,1]$ and exhibits pronounced oscillations with neutrino energy.
@@ C. Concurrence Fill
Figure 5: Concurrence $Q$ (red), concurrence fill $C_{F}$ (blue), and concurrence triangles illustrating the energy dependence of bipartite and tripartite entanglement in the three flavor neutrino system. The left panel shows $Q$ and $C_{F}$ in vacuum (dashed curve) and matter (solid curve), with vertical markers indicating the first, intermediate, and second oscillation maxima. Matter effects modify the distribution of quantum correlations, leading to characteristic enhancements or suppressions relative to vacuum. The right panel displays concurrence triangles constructed from $(C_{e\mu},C_{e\tau},C_{\mu\tau})$, where the enclosed area provides a geometric indicator of W type tripartite entanglement. Together, the panels summarize the dynamical redistribution of entanglement across energies relevant to long-baseline experiments such as DUNE.
From the analysis of the partial tangle and concurrence, we find compelling evidence of bipartite entanglement embedded within the tripartite structure of neutrino oscillations. Furthermore, the behaviour of the total tangle indicates that the three-flavor neutrino system effectively exhibits the characteristics of a W-type entangled state. To achieve a more quantitative and geometrically intuitive characterisation of this behaviour, we now turn to the study of the concurrence fill, $C_{F}$, which serves as a robust and faithful geometric measure of the underlying entanglement structure of a qutrit system.
@@
Table 5: The table lists the survival and appearance probabilities ($P_{\mu\mu},P_{\mu e},P_{\mu\tau}$) together with the three bipartite concurrences ($C^{2}_{\mu|e\tau}$, $C^{2}_{e|\mu\tau}$, $C^{2}_{\tau|\mu e}$), concurrence triangle perimeter $Q$, the concurrence fill $C_{F}$ capturing genuine tripartite (W-type) correlations, and the parameter $k$ entering the W-class inequality. These values illustrate how the pattern of bipartite and tripartite entanglement varies with energy: strong appearance probabilities correlate with enhanced bipartite concurrences, while nonzero $C_{F}$ identifies energy regions where genuine tripartite entanglement arises. The listed energies correspond to characteristic points in the oscillation spectrum, highlighting the dynamical redistribution of quantum correlations across different oscillation regimes.
@@
To explore this behaviour in detail, we considere four benchmark values energy as mentioned before. The corresponding concurrence triangles are shown in the left panel of Fig. 5. At the first and second oscillation maxima, the area of the triangle is minimal, with $C_{F}=0.06$. At these energies, the survival probability $P_{\mu\mu}$ is nearly zero, while the appearance probabilities $P_{\mu e}$ and $P_{\mu\tau}$ reach their maxima (Fig.2,Table 3). Consequently, the entanglement between the $\mu$-$\tau$ flavor modes is strongly suppressed (Fig. 4), represented by the shorter sides of the triangle (arms $A$ and $B$). In contrast, the entanglement between the appearance channels $e$-$\tau$ becomes dominant. The equal area of the concurrence triangles at these two energy points ($C_{F}=0.06$) indicates that the degree of entanglement shared between the $e$ and $\tau$ flavors is comparable in both cases. At $E=1.3~\mathrm{GeV}$, the concurrence fill $C_{F}$ vanishes entirely, corresponding to a situation where the concurrence triangle collapses into a straight line. This signifies the absence of genuine tripartite entanglement, marking a transition to a biseparable configuration. The area reaches its maximum value, $C_{F}\simeq 0.4$, at $E=1.72~\mathrm{GeV}$, where both appearance and disappearance probabilities are nonzero. At this energy, the probabilities $P_{\mu\mu}$ and $P_{\mu\tau}$ become nearly equal, leading to two sides of comparable length, $C_{\mu|e\tau}$ and $C_{\tau|e\mu}$. The third side, associated with $C_{e|\mu\tau}$, also contributes significantly due to the finite value of $P_{\mu e}$. A quantitative comparison of the triangle areas, summarised in Table 5, further substantiates this observation. The values $C_{F}(2.6~\mathrm{GeV})=C_{F}(0.8~\mathrm{GeV})=0.06$ correspond to the minima, while $C_{F}(1.7~\mathrm{GeV})=0.4$ marks the maximum. Thus, $C_{F}(1.7~\mathrm{GeV})>C_{F}(0.8~\mathrm{GeV}),C_{F}(2.5~\mathrm{GeV})$, indicating that the entanglement strength peaks at 1.7 GeV, where the interplay among all three flavor modes is most pronounced. The larger value of $C_{F}$ at this energy reflects a higher degree of collective flavor-mode genuine tripartite entanglement within the neutrino system.
The nonzero area of the concurrence triangle serves as a direct indicator of genuine tripartite entanglement in the three-flavor neutrino system. However, the mere presence of tripartite entanglement does not specify its class. The underlying state may belong either to the W family or to the GHZ family, which represent fundamentally different modes of multipartite correlation. To determine which class is realized during neutrino oscillations, it is essential to apply a more discriminating criterion. For this purpose, we now turn to an inequality-based analysis that provides a quantitative method for distinguishing W-type entanglement from GHZ-type or biseparable configurations.
@@
From the above analysis, we clearly see that the three-flavor neutrino system possesses a genuine tripartite entanglement, which can in principle belong either to the W-type or GHZ-type class of states. To identify which class the neutrino system corresponds to, we make use of the following inequality, which serves as a distinguishing criterion for W-type entanglement:
@@
At $E\simeq 1.7~\mathrm{GeV}$, where the concurrence area is largest, the inequality is strongly satisfied, signalling a robust W-type entanglement among all three flavor modes. Interestingly, this condition continues to hold over a broad energy window, roughly between $3$ and $8~\mathrm{GeV}$, suggesting that higher-energy neutrinos tend to maintain W-type correlations more dominantly. Overall, the inequality analysis provides a clear and quantitative confirmation that the entanglement structure in the three-flavor neutrino system belongs primarily to the W class, rather than the GHZ class, of tripartite states.
@@ 5 Conclusion
In the present study, we have elaborately examined the entangled structure of three flavor neutrino oscillations within a three qubit framework, highlighting that a single neutrino, although produced in a definite flavor state, naturally evolves into a coherent excitation distributed across three orthogonal flavor modes. This viewpoint shows that neutrino oscillations encompass not only quantum superposition among mass eigenstates but also constitute a genuine single particle multimode entanglement phenomenon in the flavor basis.
We developed a comprehensive formalism to quantify this multimode structure through three complementary tripartite entanglement measures: (i) the tangle, (ii) the partial tangle, and (iii) the concurrence fill. Importantly, each of these entanglement measures are expressed in terms of experimentally accessible appearance and disappearance probabilities. Using the density-matrix formalism including the matter effects, we derived analytic expressions for bipartite and bi-partitioned concurrences, the partial tangles, the three-tangle, and the concurrence fill. The ability to express all entanglement measures directly in terms of observable oscillation probabilities provides a bridge between quantum information theory and long baseline neutrino experiments. In particular, experiments such as DUNE, with broad energy coverage and high precision, provide an excellent platform for probing the dynamical behavior and distribution of multimode entanglement in the neutrino sector.
The DUNE experiment provides an exceptionally favorable setting for quantum informational studies of neutrino oscillations, owing to its broad band neutrino energy spectrum and substantial flux across the $0.5-3~\mathrm{GeV}$ range. This wide coverage allows DUNE to probe not only the first oscillation maximum but also the second where the oscillation amplitude of $P_{\mu e}$ is amplified by a factor of three as compared to the first oscillation. The ability to access both oscillation maxima is crucial for resolving the dynamical redistribution of flavor mode entanglement in ways that narrow band beams cannot achieve. Motivated by these advantages, the present work adopted the DUNE experimental configuration as the primary framework for our quantum informational analysis. All numerical results have been obtained using detailed simulations performed with the GLoBES software package, ensuring that the behavior of the entanglement measures reflects realistic long baseline oscillation dynamics.
@@
The partial-tangle analysis further clarifies the monogamy structure of neutrino entanglement. The three partial tangles exhibit clear oscillatory and anti-correlated patterns: close to the first oscillation maximum, one partial tangle, $\tau_{e\tau}$, becomes nearly maximal while the others, $\tau_{\mu\tau}$, $\tau_{\mu e}$, are strongly suppressed. This behavior illustrates explicit monogamy of entanglement i.e., enhancement of correlations in one pair ($\tau_{e\tau}$) leads to diminishing correlations in the remaining pairs ($\tau_{\mu\tau}$, $\tau_{\mu e}$). The redistribution of partial tangles preserves the topology of W-type entanglement with vanishing tangle for each flavor.
@@
We then computed the concurrence fill $C_{F}$ which is a geometric measure of genuine tripartite entanglement for the three flavors neutrino oscillations. The concurrence fill reflects the area enclosed by the concurrence triangle with each side representing the bi-partitioned concurrence $C_{\alpha|\beta\gamma}$. The estimated area of the concurrence triangle varies with energy and it becomes minimum. The non-vanishing values for $C_{F}$ over most energies manifests the presence of W-type entanglement during the flavor evolution. Further, W-class inequality provides a decisive classification of the entanglement structure. It is observed that, at $E=1.3$ GeV, all the bipartite and tripartite entanglement measures vanish identically and the neutrino system behaves completely separable.
@@
In summary, the present analysis demonstrates that the dynamical evolution of a three-flavor neutrino system provide a natural realization of single-particle multimode entanglement with its structure fully expressible in terms of experimentally measurable flavor transition oscillation probabilities. The framework developed here enables an unified method for identifying separable, biseparable, and W-type configurations as a functions of energy and baseline, thereby establishing a direct bridge between quantum information diagnostics and long baseline neutrino phenomenology. It is important to investigate how these entanglement measures behave in the presence of nonstandard interactions, decay or decoherence effects Dixit:2020ize ; Chattopadhyay:2021eba and can impact the precision measurements of oscillation parameters. The present formalism can be extended naturally to scenarios with sterile neutrinos, nonunitary mixing, or in a events of astrophysical origin i.e., in supernovae scenarios Dighe:1999bi ; Das:2017iuj ; Dighe:2017sur ; Sen:2024fxa . Thus, with the next generation facilities such as DUNE, Hyper-K, and ESSnuSB entering high precision regimes, the entanglement-based tools provided here offer a promising avenue for probing quantum-information aspects of neutrino flavor evolution in experimentally accessible settings.

### v2

@@ C. Tangle (or Three-Tangle) and Partial Tangle
Figure 4: Energy dependence of the partial tangles $\tau^{2}_{i|jk}$ in the three-flavor neutrino system for vacuum (dashed) and matter (solid) propagation. The curves correspond to the three bipartitions $\tau^{2}_{\mu e}$ (tuquoise), $\tau^{2}_{e\tau}$ (purple), and $\tau^{2}_{\mu\tau}$ (yellow), and quantify how the residual entanglement associated with each flavor mode is shared among the remaining two modes. Vertical markers indicate representative energies (2.6, 1.7, 1.3, and 0.8 GeV) corresponding to the first, intermediate, and second oscillation maxima. The variation of the partial tangles demonstrates that the distribution of residual entanglement evolves non-uniformly with energy. In vacuum, the structure reflects the coherent interplay among flavor components, whereas matter effects further reshape this pattern by enhancing or suppressing individual bipartitions. Overall, the figure provides a compact view of how tripartite entanglement dynamically reorganizes as the neutrino propagates, offering insight into the quantum structure relevant to long-baseline experiments such as DUNE.
@@
We next examine the partial tangle, $\tau_{\alpha\beta}$, which probes the residual bipartite entanglement between a selected pair $(\alpha,\beta)$ within a correlated tripartite system. Figure 4 displays the energy dependence of the three partial tangles. $\{\tau_{\mu e}^{2}\ \text{(cyan)},\ \tau_{\mu\tau}^{2}\ \text{(yellow)},\ \tau_{e\tau}^{2}\ \text{(purple)}\}$ in both vacuum (dashed) and matter (solid) across the entire accessible energy range of DUNE. We observe that each partial tangle lies in the interval $[0,1]$ and exhibits pronounced oscillations with neutrino energy.
@@ C. Concurrence Fill
Figure 5: Concurrence $Q$ (red), concurrence fill $C_{F}$ (blue), and concurrence triangles illustrating the energy dependence of bipartite and tripartite entanglement in the three flavor neutrino system. The left panel shows $Q$ and $C_{F}$ in the vacuum (dashed curves) and in matter (solid curves), with vertical markers indicating the first, intermediate, and second oscillation maxima. Matter effects modify the distribution of quantum correlations, leading to characteristic enhancements or suppressions relative to the vacuum. The right panel displays concurrence triangles constructed from $(C_{e\mu},C_{e\tau},C_{\mu\tau})$, where the enclosed area provides a geometric indicator of W type tripartite entanglement. Together, the panels summarize the dynamical redistribution of entanglement across energies relevant to long-baseline experiments such as DUNE.
From the analysis of the partial tangle and concurrence, we find compelling evidence of bipartite entanglement embedded within the tripartite structure of neutrino oscillations. To achieve a more quantitative and geometrically intuitive characterisation of this behaviour, we now turn to the study of the concurrence fill, $C_{F}$, which serves as a robust and faithful geometric measure of the underlying entanglement structure of a qutrit system.
@@
Table 5: The table lists the survival and appearance probabilities ($P_{\mu\mu},P_{\mu e},P_{\mu\tau}$) together with the three bipartite concurrences ($C^{2}_{\mu|e\tau}$, $C^{2}_{e|\mu\tau}$, $C^{2}_{\tau|\mu e}$), concurrence triangle perimeter $Q$, the concurrence fill $C_{F}$ capturing genuine tripartite (W-type) correlations. The parameter $k$ enters the W-class inequality. These values illustrate how the pattern of bipartite and tripartite entanglement varies with energy: strong appearance probabilities correlate with enhanced bipartite concurrences, while nonzero $C_{F}$ identifies energy regions where genuine tripartite entanglement arises. The listed energies correspond to characteristic points in the oscillation spectrum, highlighting the dynamical redistribution of quantum correlations across different oscillation regimes.
@@
To explore this behaviour in detail, we consider four benchmark energy values, as mentioned before. The corresponding concurrence triangles are shown in the left panel of Fig. 5. At the first and second oscillation maxima, the area of the triangle is minimal, with $C_{F}=0.06$. At these energies, the survival probability $P_{\mu\mu}$ is nearly zero, while the appearance probabilities $P_{\mu e}$ and $P_{\mu\tau}$ reach their maxima (Fig.2,Table 3). Consequently, the entanglement between the $\mu$-$\tau$ flavor modes is strongly suppressed (Fig. 4), represented by the shorter sides of the triangle (arms $A$ and $B$). In contrast, the entanglement between the appearance channels $e$-$\tau$ becomes dominant. The equal area of the concurrence triangles at these two energy points ($C_{F}=0.06$) indicates that the degree of entanglement shared between the $e$ and $\tau$ flavors is comparable in both cases. At $E=1.3~\mathrm{GeV}$, the concurrence fill $C_{F}$ vanishes entirely, corresponding to a situation where the concurrence triangle collapses into a straight line. This signifies the absence of genuine tripartite entanglement, marking a transition to a biseparable configuration. The area reaches its maximum value, $C_{F}\simeq 0.4$, at $E=1.72~\mathrm{GeV}$, where both appearance and disappearance probabilities are nonzero. At this energy, the probabilities $P_{\mu\mu}$ and $P_{\mu\tau}$ become nearly equal, leading to two sides of comparable length, $C_{\mu|e\tau}$ and $C_{\tau|e\mu}$. The third side, associated with $C_{e|\mu\tau}$, also contributes significantly due to the finite value of $P_{\mu e}$. A quantitative comparison of the triangle areas, summarised in Table 5, further substantiates this observation. The values $C_{F}(2.6~\mathrm{GeV})=C_{F}(0.8~\mathrm{GeV})=0.06$ correspond to the minima, while $C_{F}(1.7~\mathrm{GeV})=0.4$ marks the maximum. Thus, $C_{F}(1.7~\mathrm{GeV})>C_{F}(0.8~\mathrm{GeV}),C_{F}(2.5~\mathrm{GeV})$, indicating that the entanglement strength peaks at 1.7 GeV, where the interplay among all three flavor modes is most pronounced. The larger value of $C_{F}$ at this energy reflects a higher degree of genuine tripartite entanglement between the flavor modes within the neutrino system.
The nonzero area of the concurrence triangle serves as a direct indicator of genuine tripartite entanglement in the three-flavor neutrino system. However, the mere presence of tripartite entanglement does not specify its class. The underlying state may belong either to the W family or to the GHZ family, which represents fundamentally different modes of multipartite correlation. To determine which class is realized during neutrino oscillations, it is essential to apply a more discriminating criterion. For this purpose, we now turn to an inequality-based analysis that provides a quantitative method for distinguishing W- from GHZ-type or biseparable configurations.
@@
From the above analysis, we find that for most of the chosen benchmark energies, the neutrino state exhibits genuine tripartite entanglement. However, at one particular benchmark point, $E=1.3\,\mathrm{GeV}$, the state approaches a nearly biseparable configuration, indicating a qualitative change in the underlying entanglement structure. To identify which class the neutrino system corresponds to, we make use of the following inequality, which serves as a distinguishing criterion for W-type entanglement:
@@
At $E\simeq 1.7~\mathrm{GeV}$, where the concurrence area is largest, the inequality is strongly satisfied, signaling a robust W-type entanglement among all three flavor modes. Interestingly, this condition continues to hold over a broad energy window, roughly between $3$ and $8~\mathrm{GeV}$, suggesting that higher-energy neutrinos tend to maintain W-type correlations more dominantly. Overall, the inequality analysis provides a clear and quantitative confirmation that the entanglement structure in the three-flavor neutrino system belongs primarily to the W class, rather than the GHZ class of tripartite states.
E. Entanglement oscillograms and CP-phase sensitivity
It will be interesting to study the sensitivity of the genuine tripartite entanglement measure like concurrence fill $C_{F}$ to the Dirac CP phase $\delta_{CP}$ and compare it with the standard oscillation fit. It is also important to identify the kinematic regions where the concurrence fill is most sensitive to the $\delta_{CP}$. To further clarify the physical role of the concurrence fill $C_{F}$ in probing $\delta_{CP}$, we present oscillograms showing the variation of $C_{F}$ as a function of neutrino energy $E$ and $\delta_{CP}$ in Fig. 7. In these oscillograms, the vertical axis represents the neutrino energy $E$, the horizontal axis corresponds to the CP-violating phase $\delta_{CP}$, and the color scale encodes the magnitude of the concurrence fill $C_{F}$. Each point in the plot, therefore, corresponds to a fully evolved three-flavor neutrino state at fixed baseline, with the color indicating the strength of W-type tripartite correlations among the flavor modes. The vacuum panel illustrates purely kinematic interference among mass eigenstates, while the matter panel incorporates the additional modifications induced by the matter potential.
From the perspective of CP-phase sensitivity, the key quantity is the variation of $C_{F}$ with respect to $\delta_{CP}$ i.e., the magnitude of the gradient, $\left|\frac{\partial C_{F}}{\partial\delta_{CP}}\right|$ which geometrically corresponds to the horizontal color variation in the oscillogram. The regions where the color changes rapidly along the $\delta_{CP}$ direction indicate strong phase responsiveness, meaning that small shifts in $\delta_{CP}$ produce appreciable changes in the entanglement structure. In contrast, horizontally uniform bands correspond to $\partial C_{F}/\partial\delta_{CP}\approx 0$, where the observable $C_{F}$ becomes effectively insensitive to the CP phase.
Figure 7: Oscillograms of the concurrence fill $C_{F}$ as functions of neutrino energy $E$ and the CP phase $\delta_{\rm CP}$ for a $1300~\mathrm{km}$ baseline, shown for matter (left panel) and for vacuum (right panel). The dashed horizontal lines mark representative energies corresponding to the first ($\sim 2.6~\mathrm{GeV}$) and second ($\sim 0.8~\mathrm{GeV}$) oscillation maxima, along with intermediate benchmark points ($E\sim 1.3~\mathrm{GeV},1.7~\mathrm{GeV}$). The color scale indicates the magnitude of $C_{F}$. Comparing the two panels demonstrates that matter effects reshape the distribution of genuine tripartite entanglement across the energy spectrum, modifying both the location and the intensity of regions where $C_{F}$ is enhanced or suppressed.
Next, we examine, in this context, three interesting kinematic regions corresponding to the energies $E\simeq 1.3$ GeV (minimal mixing and minimal $C_{F}$); $E\simeq 1.7$ GeV (maximal mixing and maximal $C_{F}$), and $E\simeq 2.6(0.8)$ GeV for first (2nd) oscillation maxima with intermediate values for $C_{F}$. As may be observed in Fig. 7, near $E\simeq 1.7~\mathrm{GeV}$, the concurrence fill reaches near-maximal values, signaling strong tripartite entanglement. This enhancement is not uniform in $\delta_{\mathrm{CP}}$, but is concentrated in the interval $\delta_{\mathrm{CP}}\approx 150^{\circ}\text{--}300^{\circ}$. At $E\simeq 1.3~\mathrm{GeV}$, the concurrence fill becomes strongly suppressed, and $C_{F}$ exhibits a nearly uniform horizontal band structure. Because the value of $C_{F}$ remains largely constant regardless of $\delta_{CP}$, the derivative approaches zero $\left(\frac{\partial C_{F}}{\partial\delta_{CP}}\approx 0\right)$. Hence, in this regime, the physical state is effectively insensitive to the CP phase. At energies corresponding to the first and second oscillation maxima (e.g., $E\sim 2.6~\mathrm{GeV}$ and $0.8~\mathrm{GeV}$), the entanglement measure $C_{F}$ assumes a moderate values but exhibits pronounced horizontal gradients. In these regions, small variations in $\delta_{CP}$ produce significant changes in $C_{F}$, indicating enhanced phase sensitivity, thereby making $C_{F}$ highly phase-dependent.
In particular, the regions near the oscillation maxima — especially around the second oscillation maximum where matter effects are comparatively moderate — exhibit pronounced horizontal gradients, identifying favorable energy windows for enhanced $\delta_{CP}$ resolution. Thus, such an oscillogram serves as a useful guide for identifying optimal operating kinematic regimes in long-baseline experiments. However, when the primary objective is the precision determination of $\delta_{CP}$, the experimental focus should shift from regions of maximal $C_{F}$ to those where the observable exhibits the strongest phase sensitivity. Hence, considering $C_{F}$ as a potential CP-sensitivity observable, this moderately entangled region at $E\sim 0.8\,\mathrm{GeV}$, provides the most favorable conditions for precision measurement of $\delta_{CP}$.
@@ 5 Conclusion
In the present study, we have thoroughly examined the entangled structure of three-flavor neutrino oscillations within a three-qubit framework, highlighting that a single neutrino, although produced in a definite flavor state, naturally evolves into a coherent excitation distributed across three orthogonal flavor modes. This viewpoint demonstrates that neutrino oscillations encompass not only quantum superposition among mass eigen states but also constitute a genuine single-particle multi-mode entanglement phenomenon in the flavor basis.
We developed a comprehensive formalism to quantify this multi-mode structure through three complementary tripartite entanglement measures: (i) the tangle, (ii) the partial tangle, and (iii) the concurrence fill. Notably, each of these entanglement measures is expressed in terms of experimentally accessible appearance and disappearance probabilities. Using the density-matrix formalism, including matter effects, we derived analytic expressions for bipartite and bipartitioned concurrences, partial tangles, the three-tangle, and the concurrence fill. The ability to express all entanglement measures directly in terms of observable oscillation probabilities provides a bridge between quantum information theory and long baseline neutrino experiments. In particular, experiments such as DUNE, with broad energy coverage and high precision, provide an excellent platform for probing the dynamical behavior and distribution of multi-mode entanglement in the neutrino sector.
The DUNE experiment provides an exceptionally favorable setting for quantum informational studies of neutrino oscillations, owing to its broad band neutrino energy spectrum and substantial flux across the $0.5-3~\mathrm{GeV}$ range. This wide coverage allows DUNE to probe not only the first oscillation maximum but also the second, where the oscillation amplitude of $P_{\mu e}$ is amplified by a factor of three as compared to the first oscillation. The ability to access both oscillation maxima is crucial for resolving the dynamical redistribution of flavor mode entanglement in ways that narrow band beams cannot achieve. Motivated by these advantages, the present work adopted the DUNE experimental configuration as the primary framework for our quantum informational analysis. All numerical results have been obtained from detailed simulations using the GLoBES software package, ensuring that the behavior of the entanglement measures accurately reflects realistic long-baseline oscillation dynamics.
@@
The partial-tangle analysis further clarifies the monogamy structure of neutrino entanglement. The three partial tangles exhibit clear oscillatory and anti-correlated patterns: close to the first oscillation maximum, one partial tangle, $\tau_{e\tau}$, becomes nearly maximal while the others, $\tau_{\mu\tau}$, $\tau_{\mu e}$, are strongly suppressed. This behavior illustrates explicit monogamy of entanglement, i.e., the enhancement of correlations in one pair ($\tau_{e\tau}$) leads to a diminishing of correlations in the remaining pairs ($\tau_{\mu\tau}$, $\tau_{\mu e}$). The redistribution of partial tangles preserves the topology of W-type entanglement with vanishing tangle for each flavor.
@@
We then computed the concurrence fill $C_{F}$ which is a geometric measure of genuine tripartite entanglement for the three flavors neutrino oscillations. The concurrence fill reflects the area enclosed by the concurrence triangle, with each side representing the bi-partitioned concurrence $C_{\alpha|\beta\gamma}$. The estimated area of the concurrence triangle varies with energy, reaching a minimum. The non-vanishing values for $C_{F}$ over most energies manifest the presence of W-type entanglement during the flavor evolution. Further, W-class inequality provides a decisive classification of the entanglement structure. It is observed that, at $E=1.3$ GeV, all the bipartite and tripartite entanglement measures vanish identically and the neutrino system behaves completely separable.
We have also studied the sensitivity of the genuine tripartite entanglement measure, like concurrence fill $C_{F}$, to the Dirac CP phase $\delta_{CP}$ by the gradient $\left|\partial C_{F}/\partial\delta_{CP}\right|$ for different energies within the DUNE setup. In particular, we have examined three interesting cases corresponding to the energies $E\simeq 1.3$ GeV (minimal mixing and minimal $C_{F}$); $E\simeq 1.7$ GeV (maximal mixing and maximal $C_{F}$), and $E\simeq 2.6(0.8)$ GeV for first (2nd) oscillation maxima and intermediate values for $C_{F}$. It turns out that at $E\simeq 1.3~\mathrm{GeV}$, $C_{F}$ is almost independent of $\delta_{\mathrm{CP}}$, and the gradient $\left|\partial C_{F}/\partial\delta_{CP}\right|$ is nearly zero. In this region, the tripartite entanglement measure $C_{F}$ is effectively blind to the CP phase. Analogously, at $E\simeq 1.7~\mathrm{GeV}$, the concurrence fill approaches its maximum, indicating strong tripartite entanglement, predominantly within $\delta_{\mathrm{CP}}\approx 150^{\circ}\text{--}300^{\circ}$. In contrast, near the first (e.g., $E\sim 2.6~\mathrm{GeV}$) and especially the second oscillation maxima ( $E\sim 0.8~\mathrm{GeV}$), the concurrence fill $C_{F}$ takes moderate values while the small variation in $\delta_{\mathrm{CP}}$ produces significant changes in $C_{F}$. Thus, the gradient in $C_{F}$ with respect to $\delta_{\rm CP}$ might play an important role in long baseline experiments like DUNE, identifying the interesting kinematic regions.
@@
In summary, the present analysis demonstrates that the dynamical evolution of a three-flavor neutrino system provides a natural realization of single-particle multimode entanglement with its structure fully expressible in terms of experimentally measurable flavor transition oscillation probabilities. The framework developed here provides a unified method for identifying separable, biseparable, and W-type configurations as functions of energy and baseline, thereby establishing a direct bridge between quantum information diagnostics and long baseline neutrino phenomenology. It is important to investigate how these entanglement measures behave in the presence of nonstandard interactions, decay, or decoherence effects Dixit:2020ize ; Chattopadhyay:2021eba , which can impact the precision of measurements of oscillation parameters. The present formalism can be naturally extended to scenarios with sterile neutrinos, nonunitary mixing, or astrophysical events, i.e., supernova scenarios Dighe:1999bi ; Das:2017iuj ; Dighe:2017sur ; Sen:2024fxa . Thus, with the next generation facilities such as DUNE, Hyper-K, and ESSnuSB entering high precision regimes, the entanglement-based tools provided here offer a promising avenue for probing quantum-information aspects of neutrino flavor evolution in experimentally accessible settings.